# Qipeng Kuang Ondřej Kuželka Yuanhong Wang Yuyi Wang

### v1

@@ Abstract
The Weighted First-Order Model Counting Problem (WFOMC) asks to compute the weighted sum of models of a given first-order logic sentence over a given domain. It can be solved in time polynomial in the domain size for sentences from the two-variable fragment with counting quantifiers, known as $\text{C}^{2}$. This polynomial-time complexity is also retained when extending $\text{C}^{2}$ by one of the following axioms: linear order axiom, tree axiom, forest axiom, directed acyclic graph axiom or connectedness axiom. An interesting question remains as to which other axioms can be added to the first-order sentences in this way. We provide a new perspective on this problem by associating WFOMC with graph polynomials. Using WFOMC, we define Weak Connectedness Polynomial and Strong Connectedness Polynomials for first-order logic sentences. It turns out that these polynomials have the following interesting properties. First, they can be computed in polynomial time in the domain size for sentences from $\text{C}^{2}$. Second, we can use them to solve WFOMC with all of the existing axioms known to be tractable as well as with new ones such as bipartiteness, strong connectedness, being a spanning subgraph, having $k$ connected components, etc. Third, the well-known Tutte polynomial can be recovered as a special case of the Weak Connectedness Polynomial, and the Strict and Non-Strict Directed Chromatic Polynomials can be recovered from the Strong Connectedness Polynomials, which allows us to show that these important graph polynomials can be computed in time polynomial in the number of vertices for any graph that can be encoded by a fixed $\text{C}^{2}$ sentence and a conjunction of an arbitrary number of ground unary literals.
@@ 1 Introduction
Given a first-order logic sentence $\Psi$ and $n\in\mathbb{N}$, the first-order model counting (FOMC) problem asks to compute the number of models of $\Psi$ over the domain $\{1,\dots,n\}$. FOMC can be used, among others, to solve various combinatorics problems on labeled structures. For instance, the FOMC of the sentence $\Psi=\forall x\exists y(E(x,y)\lor E(y,x))\land\forall x\lnot E(x,x)$ corresponds to the number of $n$-vertex directed graphs with no isolated vertices and no loops. Similarly, the FOMC of $\Phi=\forall x\lnot E(x,x)\land\forall x\forall y\left(E(x,y)\to E(y,x)\right)\land\forall x\exists_{=k}yE(x,y)$ equals the number of $k$-regular graphs on $n$ vertices where $\exists_{=k}y$ is a counting quantifier restricting the number of $y$ satisfying $E(x,y)=\top$ to exactly $k$ for each $x$.
@@
The weighted variant of the FOMC problem, known as the weighted first-order model counting (WFOMC) problem, additionally expects a specification of positive and negative weights for predicates in the language, which are then used to assign weights to models. In WFOMC, the task is to compute the sum of the weights of the given sentence’s models. [^1] Surprisingly, WFOMC is also essential for the unweighted FOMC problem since, for example, Skolemization for WFOMC [dBMD14], which is used to eliminate existential quantifiers, relies on the usage of (negative) weights. WFOMC has applications in Statistical Relational Learning (SRL) [GT07]. For instance, Markov Logic Networks (MLN) [RD06], one of the most popular SRL formalisms, model relational data by a distribution over possible worlds, where the computation of the partition function (equivalent to probabilistic inference in MLN) is reducible to WFOMC [dB11].
@@
A key issue regarding WFOMC is its computational complexity. Generally, there is no hope of devising algorithms for WFOMC that would scale polynomially with the size of the first-order logic sentences under reasonable assumptions, since it can be reduced from #SAT. This is true even for very simple fragments such as the one-variable fragment $\text{FO}^{1}$, which contains sentences with at most one logical variable. Therefore, most of the focus in this area has been on identifying fragments containing first-order logic sentences for which WFOMC can be computed in time polynomial in the size of the domain (but not necessarily in the size of the sentence). The term coined for this kind of tractability by [dB11] is domain liftability—it is an analog to data complexity [Var82] from database theory. Domain liftability also means that enumerative combinatorial problems can be solved in time polynomial in the size of the structures being counted.
@@
The first positive results regarding the tractability of WFOMC came from the two seminal papers [dB11, dBMD14], which together established domain liftability of the two-variable fragment of first-order logic $\text{FO}^{2}$. This was quickly complemented by a hardness result [BdBGS15], showing that WFOMC for the three-variable fragment $\text{FO}^{3}$ is not domain-liftable (under plausible complexity-theoretic assumptions). However, this does not mean that the frontiers of tractability cannot be pushed beyond $\text{FO}^{2}$. For instance, the two-variable fragment with counting quantifiers $\text{C}^{2}$ [GOR97] (e.g., the above sentence of a $k$-regular graph) with cardinality constraints (i.e., the restrictions to the number of true ground atoms for relations), which is strictly more expressive than $\text{FO}^{2}$, was shown to be domain-liftable in [Kuz21]. Other works [dBD12, WPWK24] proved domain-liftability when attaching unary evidence (i.e., a series of ground unary literals) to these fragments.
@@
One strategy for discovering new tractable fragments is to add extra axioms, which may or may not be finitely expressible in first-order logic. This approach was first explored in [KL18] where it was shown that adding a single functionality axiom to $\text{FO}^{2}$ results in a domain-liftable fragment. [^2] This fragment is contained in $\text{C}^{2}$, which can encode an arbitrary number of functionality constraints, but at that time $\text{C}^{2}$ was not yet known to be domain-liftable. Subsequent works extended domain liftability to the fragments $\text{C}^{2}$ + Tree [vBK23] and $\text{C}^{2}$ + LinearOrder [TK23], which are obtained by adding to $\text{C}^{2}$ an axiom specifying that a distinguished relation $R$ should correspond to an undirected tree or to a linear order, respectively. Recently, the work [MBS23] added several new tractable classes: $\text{C}^{2}$ + DAG, $\text{C}^{2}$ + Connected, and $\text{C}^{2}$ + Forest, which allow restricting a distinguished binary relation to be a directed acyclic graph or to be a connected undirected graph or to be a forest. Except for the linear-order and the functionality axiom, these axioms are not finitely expressible in first-order logic.
@@
What axioms can be attached to the first-order logic fragments while preserving domain-liftability?
One of the potential solutions is the axioms representing combinatorial structures derived from graph polynomials such as Tutte polynomial [Tut54] or directed chromatic polynomials [Sta70].
@@ 1.1 Our Contribution
In this paper, we take a different route inspired by graph polynomials. We consider $\text{C}^{2}$ possibly extended by cardinality constraints and unary evidence, denoted by extended $\text{C}^{2}$. We introduce several polynomials that we associate to first-order logic sentences and show that they can be computed efficiently for extended $\text{C}^{2}$ sentences, which extends the domain-liftability to WFOMC for several new axioms (as well as for all of the old ones) and enables the efficient computation of the Tutte polynomial for any graph that can be encoded by certain first-order fragments.
Polynomials for WFOMC
We first define the Weak Connectedness Polynomial (WCP) as a univariate polynomial of $u$ for any first-order sentence $\Psi$ with domain size, a pair of weighting functions, and a distinguished binary relation $R$. Roughly speaking, it is the sum over the weight of each possible graph interpreting $R$ (i.e., a graph denoted by $G(R)$ obtained by regarding each $R(a,b)=\top$ as an edge $a\to b$) multiplied by a term of $u$ indicating the number of its weakly connected components. The WCP can be regarded as a classification of all possible $G(R)$ by the number of connected components and, therefore, captures the information of the connectedness of models.
We then incorporate edge orientations by defining two Strong Connectedness Polynomials (SCPs), namely Non-strict Strong Connectedness Polynomial (NSCP) and Strict Strong Connectedness Polynomial (SSCP), which are intuitively the sum over the weight of each possible $G(R)$ multiplied by the non-strict and strict directed chromatic polynomials respectively, so that they embrace the operations of directed chromatic polynomials. As an extension of WCP, we also define Multi-relation Weak Connectedness Polynomial (MWCP) for multiple distinguished binary relations under the conditions that the relations are nested.
The polynomials are formally defined in the language of first-order logic such that the evaluation of the polynomials at each positive integer point is equivalent to the WFOMC of a certain sentence. Therefore a natural way to obtain the polynomials is to evaluate the polynomials at a sufficiently large number of points and recover the polynomials by interpolation. However, existing algorithms for WFOMC can not solve these resulting WFOMC problems in polynomial time data complexity, as in our case, the length of the sentences depends on the domain size $n$ and the runtime of these algorithms can be exponential in the length of the sentence.
We further adapt the existing framework for WFOMC of $\text{FO}^{2}$ sentences to dynamic programming algorithms involving insight into the structure of our WFOMC queries, hence we show that all these polynomials can be computed in polynomial time in the domain size.
Proposition 1.1 (Summary of Propositions 3.3, 4.3, 4.4 and 6.1).
Computing WCP, NSCP and SSCP for a given extended $\text{C}^{2}$ sentence, a domain of size $n$, a distinguished binary relation, and weighting functions can be computed in polynomial time in $n$.
Moreover, computing MWCP for a given extended $\text{C}^{2}$ sentence, a domain of size $n$, a sequence of distinguished nested binary relations, and weighting functions can be computed in polynomial time in $n$.
Lifted Counting Tasks
With the help of polynomials for WFOMC, we can tackle several counting tasks efficiently.
The first task is WFOMC for an extended $\text{C}^{2}$ sentence associated with axioms, the special constraints on $G(R)$ that it should be a certain combinatorial structure. It can be generalized to multiple relations $R_{1},\cdots,R_{k}$ such that $G(R_{1}),\cdots,G(R_{k})$ satisfy additional internal restrictions. As the polynomials allow us to obtain knowledge of combinatorial structures using similar operations as graph polynomials, this gives rise to the domain-liftability of new axioms (such as the bipartite axiom), generalization of the existing axiom (such as $connected_{k}$ from the connected axiom), and incorporation of the existing axioms (such as the tree axiom). The main axioms for which we prove domain-liftability are listed in Table 1.
Table 1: Axioms that we give domain-liftability to. Note that the first 5 axioms require the distinguished relations to be symmetric and irreflexive.
@@
$forest(R)$ $G(R)$ is a forest Theorem 3.4 [MBS23]
$SS(R_{1},R_{2})$ $G(R_{1})$ is a spanning subgraph of $G(R_{2})$; $G(R_{2})$ has the same number of connected components with $G(R_{1})$ Proposition 6.2 new
@@
The fragment of extended $\text{C}^{2}$ with any single axiom in Table 1 is domain-liftable.
Moreover, we can give domain-liftability to the combination of axioms in Table 1. Table 2 offers a comprehensive list of potential axioms by combining the existing axioms. All of them are new axioms that have not been shown to be domain-liftable before. A common challenge in prior research was the individual treatment of axioms using specific techniques, which made their combination a non-trivial task. However, our methodology seamlessly integrates these axioms through a unified polynomial that encapsulates the structural characteristics of the graphs.
@@
The fragment of extended $\text{C}^{2}$ with a single combined axiom in Table 2 is domain-liftable.
Our approach offers several distinct advantages over existing techniques for individual axioms. First, it is more general, requiring only a polynomial representation for axioms without relying on specific algorithmic techniques for each problem. Second, it enables the use of the same polynomial for multiple axioms. Third, it is more flexible, allowing easy extension to new axioms as long as they can be expressed with polynomials, for instance, requiring $G(R)$ to represent a $k$-cycle permutation in Example 3.2 or a strongly connected tournament in Example 4.1. Lastly, it establishes a connection between the complexity of graph polynomials and the domain-liftability of WFOMC.
Contribution to Graph Polynomials
Besides using the idea of graph polynomials to solve several counting tasks efficiently, conversely, the polynomials for WFOMC make contribution to graph polynomials. We show that the well-known Tutte polynomial [Tut54] can be recovered from the extended version of WCP, and the strict and non-strict directed chromatic polynomials can be recovered from SSCP and NSCP respectively, which allows us to show that these graph polynomials can be computed in time polynomial in the number of vertices for any graph that can be encoded by a fixed $\text{C}^{2}$ sentence with cardinality constraints and unary evidence.
Theorem 1.3 (Informal summary of Theorem 3.5 and Corollary 4.1).
Let $G$ be a graph that can be encoded by a $\text{C}^{2}$ sentence of constant size associated with cardinality constraints and unary evidence. The Tutte polynomial (if $G$ is undirected), the strict directed chromatic polynomial, and the non-strict directed chromatic polynomial of $G$ can be computed in time polynomial in the size of $G$.
The above result answers in the positive way the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph can be done in time polynomial in the number of vertices. However, this result cannot be extended to computing Tutte polynomial or evaluating any point for a general graph, since the first-order sentence encoding the given graph might not satisfy the condition in the theorem.
@@ 1.2 Related Work
We build on several lines of research. First, we build on the stream of results from lifted inference literature which established domain-liftability of several natural fragments of first-order logic [Poo03, dSBAR05, dB11, GD11, dBMD14, BdBGS15, KKVP16, Kuz21]. These works started with the seminal results showing domain-liftability of $\text{FO}^{2}$ [dB11, dBMD14] and continued by showing domain-liftability of larger fragments such as S2FO2 and S2RU [KKVP16], U1 [KL18] or $\text{C}^{2}$ [Kuz21]. Second, we build on extensions of these works that added axioms not expressible in these fragments or even in first-order logic such as the tree axiom [vBK23], forest axiom [MBS23], linear order axiom [TK23], acyclicity axiom [MBS23] and connectedness axiom [MBS23]. In this work, we generalize all of these results and add new axioms to this list, showing their domain-liftability. Finally, our work builds on tools from enumerative combinatorics [SS86] and on graph polynomials [Tut04]. The most well-known among graph polynomials is the Tutte polynomial [Tut54] but there exist also extensions of it [AB20], with which we do not work directly in this paper. Computing the Tutte polynomial is known to be intractable in general [JVW90], however, in this work we show that it can be computed in polynomial time for graphs that can be encoded by a fixed sentence from the first-order fragment $\text{C}^{2}$, possibly with cardinality constraints and unary evidence.
@@ 2.2 First-Order Logic
We consider the function-free, finite domain fragment of first-order logic. An atom of arity $k$ takes the form $P(x_{1},\cdots,x_{k})$ where $P/k$ is from a vocabulary of predicates (also called relations), and $x_{1},\cdots,x_{k}$ are logical variables from a vocabulary of variables.  [^3] We restrict $k\geq 1$ for simplicity but $k\geq 0$ is tractable for this work as well. A literal is an atom or its negation. A formula is formed by connecting one or more literals together using negation, conjunction, or disjunction. A formula may optionally be surrounded by one or more quantifiers of the form $\forall x$ or $\exists x$, where $x$ is a logical variable. A logical variable in a formula is said to be free if it is not bound by any quantifier. A formula with no free variables is called a sentence. As usual in the literature, the equality symbol “=” is not allowed. [^4] Though, the identity relation can be expressed by cardinality constraints as shown at the end of Section 2.4.
A first-order logic formula in which none of the atoms contains any variable is called ground. A possible world $\omega$ interprets each predicate in a sentence over a finite domain, represented as a set of true or false ground atoms. Sometimes we only care about the interpretation of a single predicate $R$ in a possible world $\omega$, which is denoted by $\omega_{R}$. The satisfaction relation $\models$ is defined in the usual way: $\omega\models\alpha$ means that the formula $\alpha$ is true in $\omega$. The world $\omega$ is a model of a sentence $\Psi$ if $\omega\models\Psi$. We denote the set of all models of a sentence $\Psi$ over the domain $\{1,2,\dots,n\}$ by $\mathcal{M}_{\Psi,n}$.

### v2

@@ Abstract
The Weighted First-Order Model Counting Problem (WFOMC) asks to compute the weighted sum of models of a given first-order logic sentence over a given domain. It can be solved in time polynomial in the domain size for sentences from the two-variable fragment with counting quantifiers, known as $\text{C}^{2}$. This polynomial-time complexity is known to be retained when extending $\text{C}^{2}$ by one of the following axioms: linear order axiom, tree axiom, forest axiom, directed acyclic graph axiom or connectedness axiom. An interesting question remains as to which other axioms can be added to the first-order sentences in this way. We provide a new perspective on this problem by associating WFOMC with graph polynomials. Using WFOMC, we define Weak Connectedness Polynomial and Strong Connectedness Polynomials for first-order logic sentences. It turns out that these polynomials have the following interesting properties. First, they can be computed in polynomial time in the domain size for sentences from $\text{C}^{2}$. Second, we can use them to solve WFOMC with all of the existing axioms known to be tractable as well as with new ones such as bipartiteness, strong connectedness, having $k$ connected components, etc. Third, the well-known Tutte polynomial can be recovered as a special case of the Weak Connectedness Polynomial, and the Strict and Non-Strict Directed Chromatic Polynomials can be recovered from the Strong Connectedness Polynomials.
@@ 1 Introduction
Given a function-free first-order logic sentence $\Psi$ and $n\in\mathbb{N}$, the first-order model counting (FOMC) problem asks to compute the number of models of $\Psi$ over the domain $\{1,\dots,n\}$. FOMC can be used, among others, to solve various combinatorics problems on labeled structures. For instance, the FOMC of the sentence $\Psi=\forall x\exists y(E(x,y)\lor E(y,x))\land\forall x\lnot E(x,x)$ corresponds to the number of $n$-vertex directed graphs with no isolated vertices and no loops. Similarly, the FOMC of $\Phi=\forall x\lnot E(x,x)\land\forall x\forall y\left(E(x,y)\to E(y,x)\right)\land\forall x\exists_{=k}yE(x,y)$ equals the number of $k$-regular graphs on $n$ vertices where $\exists_{=k}y$ is a counting quantifier restricting the number of $y$ satisfying $E(x,y)=\top$ to be exactly $k$ for each $x$.
@@
The weighted variant of the FOMC problem, known as the weighted first-order model counting (WFOMC) problem, additionally expects a specification of positive and negative weights for predicates in the language, which are then used to assign weights to models. In WFOMC, the task is to compute the sum of the weights of the given sentence’s models. [^1] Surprisingly, WFOMC is also essential for the unweighted FOMC problem since, for example, Skolemization for FOMC [dBMD14], which is used to eliminate existential quantifiers, relies on the usage of (negative) weights. WFOMC has applications in Statistical Relational Learning (SRL) [GT07]. For instance, Markov Logic Networks (MLNs) [RD06], one of the most popular SRL formalisms, model relational data by a distribution over possible worlds, where the computation of the partition function (used in probabilistic inference in MLN) is reducible to WFOMC [dB11].
@@
A key issue regarding WFOMC is its computational complexity. Generally, there is no hope of devising algorithms for WFOMC that would scale polynomially with the size of the first-order logic sentences under reasonable assumptions [BdBGS15]. This is true even for very simple fragments such as the one-variable fragment $\text{FO}^{1}$, which contains sentences with at most one logical variable. Therefore, most of the focus in this area has been on identifying fragments of first-order logic for which WFOMC can be computed in time polynomial in the size of the domain (but not necessarily in the size of the sentence). The term coined for this kind of tractability by [dB11] is domain liftability—it is an analog to data complexity [Var82] from database theory.
@@
The first positive results regarding the tractability of WFOMC came from the two seminal papers [dB11, dBMD14], which together established domain liftability of the two-variable fragment of first-order logic $\text{FO}^{2}$. This was quickly complemented by a hardness result [BdBGS15], showing that WFOMC for the three-variable fragment $\text{FO}^{3}$ is not domain-liftable (under plausible complexity-theoretic assumptions). However, this does not mean that the frontiers of tractability cannot be pushed beyond $\text{FO}^{2}$. For instance, the two-variable fragment with counting quantifiers $\text{C}^{2}$ [GOR97] (e.g., the above sentence of a $k$-regular graph) with cardinality constraints (i.e., the restrictions to the number of true ground atoms for relations), which is strictly more expressive than $\text{FO}^{2}$, was shown to be domain-liftable in [Kuz21]. Other works [dBD12, WPWK23, WPWK24] proved domain-liftability when attaching ground unary literals (as part of the input along with the domain size) to sentences from these fragments.
@@
One strategy for discovering new tractable fragments is to add extra axioms to a fragment known to be tractable, which may or may not be finitely expressible in first-order logic. This approach was first explored in [KL18] where it was shown that adding a single functionality axiom to $\text{FO}^{2}$ results in a domain-liftable fragment. [^2] This fragment is contained in $\text{C}^{2}$, which can encode an arbitrary number of functionality constraints. Subsequent works extended domain liftability to the fragments $\text{C}^{2}$ + Tree [vBK23] and $\text{C}^{2}$ + LinearOrder [TK23], which are obtained by adding to $\text{C}^{2}$ an axiom specifying that a distinguished relation $R$ should correspond to an undirected tree or to a linear order, respectively. Recently, the work [MBS23] added several new tractable classes: $\text{C}^{2}$ + DAG, $\text{C}^{2}$ + Connected, and $\text{C}^{2}$ + Forest, which allow restricting a distinguished binary relation to be a directed acyclic graph or to be a connected undirected graph or to be a forest. Except for the linear-order and the functionality axiom, these axioms are not finitely expressible in first-order logic.
@@
Can we define a general framework for adding axioms to $\text{C}^{2}$ that allows us to prove domain-liftability of new axioms and to efficiently compute the WFOMC of sentences with these axioms?
@@ 1.1 Our Contribution
In this paper, we take a route inspired by graph polynomials such as the Tutte polynomial [Tut54] and directed chromatic polynomials [Sta70]. Graph polynomials are algebraic expressions that encode various combinatorial properties of graphs, allowing us to capture information such as connectivity, colorability, and spanning substructures. In this work, we consider $\text{C}^{2}$ possibly extended by cardinality constraints and unbounded ground unary literals (which we will call $\text{C}^{2}$ with cardinality constraints for simplicity). We introduce similar polynomials for first-order logic sentences, capturing certain combinatorial properties of their models, and show that they can be computed efficiently for $\text{C}^{2}$ with cardinality constraints, which requires introduction of new algorithmic techniques for computation of WFOMC. This, in turn, allows us to extend the domain-liftability to WFOMC for several new axioms (as well as for all of the old ones) and enables the efficient computation of graph polynomials for graphs that can be encoded by certain first-order fragments.
Using the new polynomials for WFOMC, we can tackle the problem of efficiently computing WFOMC of $\text{C}^{2}$ sentences with cardinality constraints and one of a number of non-trivial axioms, special constraints on the graph interpreting a distinguished binary relation $R$ (denoted by $G(R)$). The main axioms for which we prove domain-liftability are listed in Table 1.
Table 1: Axioms that we proved the domain-liftability of. Note that the first 4 axioms require the distinguished relations to be symmetric and irreflexive.
@@
[…]
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and any single axiom in Table 1 is domain-liftable.
Moreover, we can show domain-liftability of combinations of axioms. Table 2 offers a comprehensive list of axioms obtained by combining the existing axioms, all of which are new axioms that have not been shown to be domain-liftable before. [^3] This would be difficult to achieve using existing techniques based on individual treatment of each axiom, as done in the literature.
@@
The fragment of $\text{C}^{2}$ with cardinality constraints and any single axiom in Table 2 is domain-liftable.
Besides using the idea of graph polynomials to solve several counting tasks efficiently, conversely, the polynomials for WFOMC bring new results on graph polynomials. We show that the Tutte polynomial [Tut54], the strict and non-strict directed chromatic polynomials [Sta70] can be recovered from our polynomials for WFOMC. This allows us to show that these graph polynomials can be computed in time polynomial in the number of vertices for any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence (the parameters of cardinality constraints and the number of ground unary literals are allowed to depend on the number of vertices). This result answers in the positive way the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph (see its definition in Example 2.10) can be done in time polynomial in the number of vertices.
@@ 1.2 Related Work
We build on several lines of research. First, we build on the stream of results from lifted inference literature which established domain-liftability of several natural fragments of first-order logic [Poo03, dSBAR05, dB11, GD11, dBMD14, BdBGS15, KKVP16, Kuz21]. These works started with the seminal results showing domain-liftability of $\text{FO}^{2}$ [dB11, dBMD14] and continued by showing domain-liftability of larger fragments such as S2FO2 and S2RU [KKVP16], U1 [KL18] or $\text{C}^{2}$ [Kuz21]. Second, we build on extensions of these works that added axioms not expressible in these fragments or even in first-order logic such as the tree axiom [vBK23], forest axiom [MBS23], linear order axiom [TK23], acyclicity axiom [MBS23] and connectedness axiom [MBS23]. In this work, we generalize all of these results and add new axioms to this list, showing their domain-liftability. Finally, our work builds on tools from enumerative combinatorics [SS86] and on graph polynomials [Tut04]. The most well-known among graph polynomials is the Tutte polynomial [Tut54] but there exist also extensions of it [AB20], with which we do not work directly in this paper. Computing the Tutte polynomial is known to be intractable in general [JVW90], however, in this work we show that it can be computed in polynomial time for graphs that can be encoded by a fixed sentence from the first-order fragment $\text{C}^{2}$, possibly with cardinality constraints and ground unary literals, both of which may depend on the number of vertices of the graph (unlike the $\text{C}^{2}$ sentence which needs to stay fixed).
@@ 2.2 First-Order Logic
We consider the function-free (we allow constants but not any functions with arity greater than zero), finite domain fragment of first-order logic. An atom of arity $k$ takes the form $P(x_{1},\cdots,x_{k})$ where $P/k$ is from a vocabulary of predicates (also called relations), and $x_{1},\cdots,x_{k}$ are logical variables from a vocabulary of variables or constants from a vocabulary of constants. [^4] We restrict $k\geq 1$ for simplicity but $k\geq 0$ is tractable for this work as well. A literal is an atom or its negation. A formula is a literal or formed by connecting one or more formulas together using negation, conjunction, or disjunction. A formula may be optionally surrounded by one or more quantifiers of the form $\forall x$ or $\exists x$, where $x$ is a logical variable. A logical variable in a formula is said to be free if it is not bound by any quantifier. A formula with no free variables is called a sentence.
A first-order logic formula in which all variables are substituted by constants in the domain is called ground. A possible world $\omega$ interprets each predicate in a sentence over a finite domain, represented by a set of ground literals. Sometimes we only care about the interpretation of a single predicate $R$ in a possible world $\omega$, which is denoted by $\omega_{R}$. The satisfaction relation $\models$ is defined in the usual way: $\omega\models\alpha$ means that the formula $\alpha$ is true in $\omega$. The possible world $\omega$ is a model of a sentence $\Psi$ if $\omega\models\Psi$. We denote the set of all models of a sentence $\Psi$ over the domain $\{1,2,\dots,n\}$ by $\mathcal{M}_{\Psi,n}$.